# Yuehao Bai

### v1

@@ 5.3 Further Reading
The exposition in Section 5.1 closely follows Negi and Wooldridge (2021) and Ma et al. (2022), although some of their expressions have been modified so that they could be more easily related to Sections 3 and 4 of this paper. The use of doubly-robust estimators for regression adjustment in randomized experiments goes back at least to the work of Tsiatis et al. (2008); Tu et al. (2023) and Rafi (2023) extend these results to general stratified experiments with finitely many strata. Bai et al. (2023b) and Cytrynbaum (2023a) consider regression adjustment in settings with “small” strata, in the sense of Section 4.2.1. Wang et al. (2023) study regression adjustment for general parameters defined by estimating equations. Recent work on design-based analyses of regression adjusted estimators can be found in Aronow and Middleton (2013), Wu and Gagnon-Bartsch (2018), Liu et al. (2021), Chang (2023), Chiang et al. (2023).
@@ 6 The Analysis of Cluster Randomized Experiments
Until now, our focus has been on experiments where treatment is assigned at the individual level and each individual’s outcome depends only on their own treatment. In this section, we consider randomized experiments where treatment is assigned at an aggregated level which we call a cluster: for example, when evaluating an educational intervention we may observe outcomes at the student-level, but assign treatment at the school-level, so that every student in the school receives the same treatment. There are two common explanations for why a researcher would consider cluster-level assignment as opposed to individual-level assignment: first, there could be logistical constraints on the experiment that require that every individual in a cluster receives the same treatment. Second, we may be concerned that there is treatment interference, i.e., the treatment statuses of individuals in a cluster may affect the outcomes of others. In Section 6.1 we introduce the framework and define some relevant analogs to the average treatment effect in this setting. In Section 6.2, we discuss inference in cluster randomized experiments under complete randomization and stratified block randomization.
@@ 6.2 Inference in Cluster Randomized Experiments
In the case of $\Delta^{\rm size}$, we consider the following cluster size weighted difference-in-means estimator:
@@
where $X_{g}$ denotes the variables used in stratification (see Bugni et al., 2022; Bai et al., 2023c, for details). This variance mirrors the variance $V^{\rm sbr}$ for individual-level stratified experiments defined in (8), and therefore similar inference procedures apply to cluster randomized experiments by modifying the individual-level outcomes to suitable cluster-level counterparts such as $\hat{Y}_{g}$.
@@ 6.3 Further Reading
The material in this section is most closely related to Bugni et al. (2022) and Bai et al. (2023c). Donner and Klar (2000) contains a textbook treatment of early work in cluster randomized experiments. Su and Ding (2021) and Wang et al. (2024) study regression adjustment for cluster randomized experiments without stratification (Bugni et al. (2022) and Bai et al. (2023c) discuss extensions of these results to settings with stratification). Other recent work on cluster randomized experiments (mostly from a design-based perspective) includes Imai et al. (2009), Schochet (2013), Middleton and Aronow (2015), Schochet et al. (2021), de Chaisemartin and Ramirez-Cuellar (2024).
@@ 7.1 Treatment Effect Heterogeneity and Quantile Treatment Effects
This article focused exclusively on estimation and inference of the unconditional average treatment effect. Recent work on the analysis of randomized experiments has studied inference for quantile treatment effects: see in particular Zhang and Zheng (2020), Jiang et al. (2021), and Jiang et al. (2023). We may also be interested in heterogeneity of the average treatment effect as a function of the observable characteristics (i.e., the conditional average treatment effect, or CATE). Although estimation and inference on the CATE has been an extremely active area of research in causal inference more broadly (see, for instance, Kennedy, 2023, for an in-depth discussion and further references), to our knowledge almost all of this work maintains the assumption that treatment assignment is independent and identically distributed across individuals; this precludes, for instance, stratified block randomization or pair matching. One exception is Zhang and Ma (2023), who present tests for treatment-covariate interactions and study their properties under general covariate-adaptive stratified randomization schemes.
@@ 7.2 Re-randomization
Re-randomization is a method of experimental assignment which, in a similar spirit to stratification, is intended to enforce “balance” in the covariate distributions between treatment and control groups. In a standard re-randomization procedure, researchers specify a balance criterion for the covariate values and then repeatedly generate assignments using a completely randomized design until an assignment is found which achieves an acceptable covariate distribution according to the balance criterion. An excellent historical summary is given in Morgan and Rubin (2012). Similarly to stratification, re-randomization leads to an increase in precision relative to complete randomization, and as a result inferences will be unnecessarily conservative unless this is taken into consideration. However, unlike with stratification, where correcting this issue usually amounts to simply modifying the standard errors, corrected inferences for re-randomization are sometimes further complicated by the fact that the limiting distribution of the difference-in-means estimator is not asymptotically normal. For an in-depth theoretical discussion of re-randomization, see Li et al. (2018), Li et al. (2020), Zhao and Ding (2021b), Lu et al. (2023), Branson et al. (2024), and Cytrynbaum (2024).
@@ 7.4 Imperfect Compliance and Attrition
Attrition arises when outcomes are not observed for some subset of the experimental units. This situation could arise, for instance, if researchers lose track of subjects in the experiment. As with imperfect compliance, if the decision for the unit to drop out of the experiment is not exogenous, but is instead determined by their unobserved characteristics, then the experiment no longer point identifies the average treatment effect. Standard resolutions to this issue include modelling the selection process (as in Heckman, 1979) or partial identification (as in Horowitz and Manski, 2000). A textbook treatment on the analysis of randomized experiments with attrition is given in Gerber and Green (2012), and DiNardo et al. (2006) review different methods for dealing with attrition within the context of the Moving to Opportunity (MTO) experiment. Ghanem et al. (2023) develop a test for attrition bias in randomized experiments. Fukumoto (2022) and Bai et al. (2024a) study the problem of attrition in the setting of matched pair experiments, and revisit common recommendations about whether or not to drop pairs with an attrited unit.
@@ 7.5 Network Experiments and Experiments with Interference
More generally, we could consider settings with complex patterns of interference, for instance, if individuals interact on a large network. Manski (2013) and Aronow and Samii (2017) develop the concept of effective treatments or exposure mappings, which summarize how a given unit’s outcome is affected by the treatments of other units. For example, in a network context, the exposure mapping may dictate that only the treatments of a unit’s direct links affect their outcome. The exposure mapping formulation of interference has had a major influence on the subsequent literature: examples include Leung (2020), Viviano (2020), Forastiere et al. (2021), Auerbach and Tabord-Meehan (2021), Munro et al. (2021), Li and Wager (2022), Leung (2022), Gao and Ding (2023), Park and Kang (2023), Viviano et al. (2023), Sävje (2024). Some recent papers that study general forms of interference without employing the formalism of exposure mappings are Wager and Xu (2021), Sävje et al. (2021), and Hu et al. (2022).
@@ 7.6 Randomization Inference
It is important to emphasize that, in both cases, finite-sample validity is only guaranteed for very specific choices of the null hypothesis. In particular, in either case considered above, if we instead consider the null hypothesis that the average treatment effect is equal to zero, then we would have neither the required group invariance property nor a “sharp” null which would guarantee finite-sample validity. However, a recent series of papers establishes that the randomization test can be asymptotically valid for these “weak” nulls, while retaining its finite-sample validity for the “sharp” null, as long as the test-statistic is constructed appropriately. See Chung and Romano (2013), Bugni et al. (2018), Bai et al. (2022), and Bai et al. (2023c), for examples in a super-population context, and Zhao and Ding (2021a), Wu and Ding (2021) for examples in a design-based context.
@@ 7.8 Response-Adaptive Designs and Bandit Experiments
Much of the recent work on adaptive designs is related to bandit problems and/or best arm identification. These are essentially policy learning problems (in the sense defined in Section 7.6) where the researcher wishes to choose a policy to maximize welfare, either for the participants in the experiment itself or for the broader population of interest. Bubeck et al. (2012) and Lattimore and Szepesvári (2020) provide comprehensive textbook introductions. Some recent work on this topic includes Russo (2016), Russo and Van Roy (2016), Agrawal and Goyal (2017), Kasy and Sautmann (2021), Adusumilli (2021), Lieber (2022), Kuang and Wager (2023), Kato et al. (2024). The analysis of treatment effect parameters in these contexts can be particularly challenging since the experiment was not necessarily designed to facilitate inference. Recent work on inference in adaptive experiments includes Bibaut et al. (2021), Hadad et al. (2021), Zhang et al. (2021), Adusumilli (2023), Chen and Andrews (2023), Hirano and Porter (2023).
@@ Appendix B Proofs of claims in main text
$\operatorname{Var}[\hat{\Delta}_{n}]$ $=\frac{S^{2}_{1}}{n_{1}}+\frac{S^{2}_{1}}{n_{0}}+\frac{S^{2}_{\Delta}}{n}+\left(\frac{1}{n}-\frac{1}{N}\right)S^{2}_{\Delta}$

### v2

@@ 5.3 Further Reading
The exposition in Section 5.1 closely follows Negi and Wooldridge (2021) and Ma et al. (2022), although some of their expressions have been modified so that they could be more easily related to Sections 3 and 4 of this paper. The use of doubly-robust estimators for regression adjustment in randomized experiments goes back at least to the work of Tsiatis et al. (2008); Tu et al. (2023) and Rafi (2023) extend these results to general stratified experiments with finitely many strata. Bai et al. (2024c) and Cytrynbaum (2023a) consider regression adjustment in settings with “small” strata, in the sense of Section 4.2.1. Wang et al. (2023a) study regression adjustment for general parameters defined by estimating equations. Recent work on design-based analyses of regression adjusted estimators can be found in Aronow and Middleton (2013), Wu and Gagnon-Bartsch (2018), Liu et al. (2021), Chang (2023), Chiang et al. (2023).
@@ 6 The Analysis of Cluster Randomized Experiments
Until now, our focus has been on experiments where treatment is assigned at the individual level and each individual’s outcome depends only on their own treatment. In this section, we consider randomized experiments where treatment is assigned at an aggregated level which we call a cluster: for example, when evaluating an educational intervention we may observe outcomes at the student-level, but assign treatment at the school-level, so that every student in the school receives the same treatment. This type of cluster-level assignment is extremely common in practice: see, for instance, Angrist and Lavy (2009); Banerjee et al. (2015); Crépon et al. (2015); Bruhn et al. (2016); Romero et al. (2020). There are two common explanations for why a researcher would consider cluster-level assignment as opposed to individual-level assignment: first, there could be logistical constraints on the experiment that require that every individual in a cluster receives the same treatment. Second, we may be concerned that there is treatment interference, i.e., the treatment statuses of individuals in a cluster may affect the outcomes of others. In Section 6.1 we introduce the framework and define some relevant analogs to the average treatment effect in this setting. In Section 6.2, we discuss inference in cluster randomized experiments under complete randomization and stratified block randomization.
@@ 6.2 Inference in Cluster Randomized Experiments
In the case of $\Delta^{\rm size}$, we consider the following cluster-size weighted difference-in-means estimator:
@@
where $X_{g}$ denotes the variables used in stratification (see Bugni et al., 2022; Bai et al., 2024d, for details). This variance mirrors the variance $V^{\rm sbr}$ for individual-level stratified experiments defined in (9), and therefore similar inference procedures apply to cluster randomized experiments by modifying the individual-level outcomes to suitable cluster-level counterparts such as $\hat{Y}_{g}$.
@@ 6.3 Further Reading
The material in this section is most closely related to Bugni et al. (2022) and Bai et al. (2024d). Donner and Klar (2000) contains a textbook treatment of early work in cluster randomized experiments. Su and Ding (2021) and Wang et al. (2024) study regression adjustment for cluster randomized experiments without stratification (Bugni et al. (2022) and Bai et al. (2024d) discuss extensions of these results to settings with stratification). Other recent work on cluster randomized experiments (mostly from a design-based perspective) includes Imai et al. (2009), Schochet (2013), Middleton and Aronow (2015), Schochet et al. (2021), de Chaisemartin and Ramirez-Cuellar (2024).
@@ 7.1 Treatment Effect Heterogeneity and Quantile Treatment Effects
This article focused exclusively on estimation and inference of the unconditional average treatment effect. Recent work on the analysis of randomized experiments has studied inference for quantile treatment effects: see in particular Zhang and Zheng (2020), Jiang et al. (2021), and Jiang et al. (2023). We may also be interested in heterogeneity of the average treatment effect as a function of the observable characteristics (i.e., the conditional average treatment effect, or CATE). Although estimation and inference on the CATE has been an extremely active area of research in causal inference more broadly (see, for instance, Kennedy, 2023, for an in-depth discussion and further references), to our knowledge almost all of this work maintains the assumption that treatment assignment is independent and identically distributed across individuals; this precludes, for instance, stratified block randomization or pair matching. Some exceptions are Zhang and Ma (2023), who present tests for treatment-covariate interactions and study their properties under general covariate-adaptive stratified randomization schemes, and Ding et al. (2019); Cytrynbaum (2025), who study parametric approximations to the CATE under complete randomization and finely stratified re-randomization designs, respectively.
@@ 7.2 Re-randomization
Re-randomization is a method of experimental assignment which, in a similar spirit to stratification, is intended to enforce “balance” in the covariate distributions between treatment and control groups. In a standard re-randomization procedure, researchers specify a balance criterion for the covariate values and then repeatedly generate assignments using a completely randomized design until an assignment is found which achieves an acceptable covariate distribution according to the balance criterion. An excellent historical summary is given in Morgan and Rubin (2012). Similarly to stratification, re-randomization leads to an increase in precision relative to complete randomization, and as a result inferences will be unnecessarily conservative unless this is taken into consideration. However, unlike with stratification, where correcting this issue usually amounts to simply modifying the standard errors, corrected inferences for re-randomization are sometimes further complicated by the fact that the limiting distribution of the difference-in-means estimator is not asymptotically normal (although normality can be restored using ex-post linear adjustment: see Li and Ding, 2020; Cytrynbaum, 2025, for details). For an in-depth theoretical discussion of re-randomization, see Li et al. (2018), Li et al. (2020), Zhao and Ding (2021b), Lu et al. (2023), Branson et al. (2024). Wang et al. (2023b) and Cytrynbaum (2025) study how to combine re-randomization with stratification.
@@ 7.4 Imperfect Compliance and Attrition
Attrition arises when outcomes are not observed for some subset of the experimental units. This situation could arise, for instance, if researchers lose track of subjects in the experiment. As with imperfect compliance, if the decision for the unit to drop out of the experiment is not exogenous, but is instead determined by their unobserved characteristics, then the experiment no longer point identifies the average treatment effect. Standard resolutions to this issue include modelling the selection process (as in Heckman, 1979) or partial identification (as in Horowitz and Manski, 2000). A textbook treatment on the analysis of randomized experiments with attrition is given in Gerber and Green (2012), and DiNardo et al. (2006) review different methods for dealing with attrition within the context of the Moving to Opportunity (MTO) experiment. Ghanem et al. (2023) develop a test for attrition bias in randomized experiments. Fukumoto (2022) and Bai et al. (2024b) study the problem of attrition in the setting of matched pair experiments, and revisit common recommendations about whether or not to drop pairs with an attrited unit.
@@ 7.5 Network Experiments and Experiments with Interference
More generally, we could consider settings with complex patterns of interference, for instance, if individuals interact on a large network. Manski (2013) and Aronow and Samii (2017) develop the concept of effective treatments or exposure mappings, which summarize how a given unit’s outcome is affected by the treatments of other units. For example, in a network context, the exposure mapping may dictate that only the treatments of a unit’s direct links affect their outcome. The exposure mapping formulation of interference has had a major influence on the subsequent literature: examples include Leung (2020), Viviano (2020), Forastiere et al. (2021), Auerbach and Tabord-Meehan (2021), Munro et al. (2021), Li and Wager (2022), Leung (2022), Gao and Ding (2023), Park and Kang (2023), Viviano et al. (2023), Sävje (2024). Some recent papers that study general forms of interference without employing the formalism of exposure mappings are Wager and Xu (2021), Sävje et al. (2021), Hu et al. (2022), and Faridani and Niehaus (2024).
@@ 7.6 Randomization Inference
It is important to emphasize that, in both cases, finite-sample validity is only guaranteed for very specific choices of the null hypothesis. In particular, in either case considered above, if we instead consider the null hypothesis that the average treatment effect is equal to zero, then we would have neither the required group invariance property nor a “sharp” null which would guarantee finite-sample validity. However, a recent series of papers establishes that the randomization test can be asymptotically valid for these “weak” nulls, while retaining its finite-sample validity for the “sharp” null, as long as the test-statistic is constructed appropriately. See Chung and Romano (2013), Bugni et al. (2018), Bai et al. (2022), and Bai et al. (2024d), for examples in a super-population context, and Zhao and Ding (2021a), Wu and Ding (2021) for examples in a design-based context.
@@ 7.8 Response-Adaptive Designs and Bandit Experiments
Much of the recent work on adaptive designs is related to bandit problems and/or best arm identification. These are essentially policy learning problems (in the sense defined in Section 7.6) where the researcher wishes to choose a policy to maximize welfare, either for the participants in the experiment itself or for the broader population of interest. Bubeck et al. (2012) and Lattimore and Szepesvári (2020) provide comprehensive textbook introductions. Some recent work on this topic includes Russo (2016), Russo and Van Roy (2016), Agrawal and Goyal (2017), Kasy and Sautmann (2021), Adusumilli (2021), Lieber (2022), Kuang and Wager (2023), Kato et al. (2024), Higbee (2024). The analysis of treatment effect parameters in these contexts can be particularly challenging since the experiment was not necessarily designed to facilitate inference. Recent work on inference in adaptive experiments includes Bibaut et al. (2021), Hadad et al. (2021), Zhang et al. (2021), Adusumilli (2023), Chen and Andrews (2023), Hirano and Porter (2023).
@@ Appendix B Proofs of claims in main text
$\operatorname{Var}[\hat{\Delta}_{n}]$ $=\frac{S^{2}_{1}}{n_{1}}+\frac{S^{2}_{1}}{n_{0}}-\frac{S^{2}_{\Delta}}{n}+\left(\frac{1}{n}-\frac{1}{N}\right)S^{2}_{\Delta}$